OpenStudy (anonymous):

(5x/4) - (3/x) =(1/4)

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

it just says solve.

OpenStudy (anonymous):

can anyone help??

OpenStudy (anonymous):

(5x/4)-(3/x)=(1/4) <=> ((5x^2)-12)/4x=(1/4) <=>[((5x^2)-12)/4x]-(1/4)=0 <=>[((5x^2)-12)*4-4x/4x]=0 <=>((20x^2)-48-4x)/4x=0 <=>(5x^2)-x-12=0 delta=sqrt(240) x1=1-sqrt(240)/10 x2=1+sqrt(240)/10 S={1-sqrt(240)/10,1+sqrt(240)/10}